OpenStudy (anonymous):

Suppose p varies directly with d, and p = 3 when d = 5. What is the value of d when p = 12?

OpenStudy (trantom):

20

OpenStudy (anonymous):

p varies directly with d p = Kd

OpenStudy (trantom):

p = 5/3d which means that 20=5/3*12

OpenStudy (anonymous):

That is the formula to for direct variation. Plug in the given to determine the constant K then do the formula again with the K and the p = 12 to solve for d

OpenStudy (trantom):

K is equal to y/x in the for y=Kx

OpenStudy (trantom):

so if p=Kd, then K= d/p

OpenStudy (trantom):

therefore you can derive that 12=(p/d)d and because K can be any instance of p/d, 3/5 works, so you just see what time 3/5 equals 12, otherwise known as multiplying 12 by 5/3